### v2

Abstract
Video Large Language Models (Video-LLMs) are flourishing and has advanced many video-language tasks. As a golden testbed, Video Question Answering (VideoQA) plays pivotal role in Video-LLM developing. This work conducts a timely and comprehensive study of Video-LLMs’ behavior in VideoQA, aiming to elucidate their success and failure modes, and provide insights towards more human-like video understanding and question answering. Our analyses demonstrate that Video-LLMs excel in VideoQA; they can correlate contextual cues and generate plausible responses to questions about varied video contents. However, models falter in handling video temporality, both in reasoning about temporal content ordering and grounding QA-relevant temporal moments. Moreover, the models behave unintuitively - they are unresponsive to adversarial video perturbations while being sensitive to simple variations of candidate answers and questions. Also, they do not necessarily generalize better. The findings demonstrate Video-LLMs’ QA capability in standard condition yet highlight their severe deficiency in robustness and interpretability, suggesting the urgent need on rationales in Video-LLM developing.
1 Introduction
Large language models (LLMs) are revolutionizing research in various X+Language tasks (with X be vision, robotics, sciences, etc.), ushering in a new era with unprecedented advancements in computer science (OpenAI, 2023; Chung et al., 2022; Touvron et al., 2023; Chiang et al., 2023; Dubey et al., 2024). Video Question Answering (VideoQA), as a multimodal and multifaceted challenge that requires both linguistic and visual reasoning, is at the forefront of this revolution. It also serves as a litmus test for evaluating the comprehensive multimodal understanding and reasoning capabilities of cutting-edge Video-LLMs 111In this paper, we denote Video-LLMs as video-language models that use LLMs with billion parameters. (Alayrac et al., 2022; Dai et al., 2023; Maaz et al., 2023; Chen et al., 2023; Liu et al., 2023; Zhang et al., 2023b; Lin et al., 2023; Li et al., 2024c; Team et al., 2023).
While VideoQA performance steadily improves with Video-LLMs, the pursuit of higher QA accuracy alone is insufficient to demonstrate the models’ capability, neither in video understanding nor in question-answering. For example, a recent study (Xiao et al., 2024) reveals that the models can answer a large portion of questions (correctly) with irrelevant video inputs or even without video at all. Also, there is a surge of works (Bai et al., 2024) studying the hallucination problem of multimodal LLMs in visual description. These observations raise the concern about whether these models can provide faithful answers in real-world applications. In this regard, we conduct a comprehensive study to Video-LLM’s behavior in VideoQA. We delve into the models’ success and failure modes from multiple aspects to provide insights towards more trustworthy and human-like machine intelligence.
We begin by collecting a series of representative Video-LLMs (Yang et al., 2022; Ko et al., 2023; Yu et al., 2023; Zhang et al., 2023a, b; Li et al., 2024c; Lin et al., 2023; Li et al., 2024a). These models have achieved state-of-the-art (SOTA) on VideoQA in either finetuned or zero-shot settings by harnessing popular LLMs such as DeBERTa-V2 (He et al., 2020), Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023), Vicuna (Chiang et al., 2023) and GPT-4 (OpenAI, 2023). The accuracy of these models reaches around 70%80% on popular benchmarks such as NExT-QA (Xiao et al., 2021) and MSVD-QA (Xu et al., 2017), for both multi-choice (MC) QA and open-ended (OE) QA.
With the models, we design a comprehensive suite of adversarial probes to dissect their behavior in either video understanding or question answering. Specifically, the probes are derived by adjusting the original VQA data or settings with targeted tests on: 1) temporal understanding, 2) visual grounding, 3) multimodal VQA reasoning, 4) robustness, and 5) generalization. More details are in Sec. 3). The analyses are thus based on a comparison of the model accuracy and specific prediction before and after the adjustment. For better comparison, we also provide the behavior of the SOTA non-LLM methods (Fu et al., 2023; Xiao et al., 2023) that fully fine-tune small language models such as BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019).
Fig. 1 presents an overview of the models’ behavior when confronted with some typical probes. It shows that the models often violate human intuition in maintaining or changing their accuracy and specific predictions. While no single model wins on all tests, we find that GPT-4o is fairly among the strongest for answering video questions though it is mainly developed for image understanding. We summarize and outline some of the common findings below, more specific analyses are presented in Sec. 3:
1.
Temporal Understanding: Most Video-LLMs answer temporal questions with high accuracy, but they can hardly reason the order of the video content and are even inferior to non-LLM methods in this regard. Notably, GPT-4o shows strong temporal reasoning capabilities.
2.
Visual Grounding: Video-LLMs significantly outperform non-LLM methods in answering video questions, but they win marginally in answer grounding. This indicates that their much better performances are largely due to their strength in capturing language priors and spurious vision-text correlations.
3.
Multimodal VQA Reasoning: Compared with non-LLM methods, Video-LLMs are better at exploiting the short-cuts in candidate answers (especially video-answer short-cuts) for multi-choice QA, reflecting their deficiency in faithful reasoning from video-question to the correct answers.
4.
Robustness: Video-LLMs are “overly robust” and unresponsive to data perturbation on videos (e.g., shuffling video frames) while being unexpectedly sensitive to language variations (e.g., rephrasing questions), especially for open-ended QA.
5.
Generalization: Fine-tuned Video-LLMs still favour the answers of high frequency as predictions in OEQA. While they generalize better across question types and datasets, this is not guaranteed for each specific model. Additionally, according to our observation on NExT-OOD (Zhang et al., 2024c), out-of-distribution (OOD) data for non-LLM methods may not be actual OOD for Video-LLMs.
On the one hand, the above findings demonstrate Video-LLM’s strengths in performing normal video question-answering. On the other hand, they highlight significant limitations in faithful visual reasoning and conditioning their predictions on the visual content. This underscores the necessity of rationales in developing future Video-LLMs.
2 VideoQA Background
2.1 Preliminaries
VideoQA is the task of answering visual content-related questions for videos. It extends image VQA (Antol et al., 2015; Goyal et al., 2017) to dynamic scenarios, such as the movement state of objects (e.g., “slow” or “fast”?, “put down” or “take away”?), action repetitions and their transitions over time (Jang et al., 2017, 2019). VideoQA is studied through multiple-choice (MCQA) and open-ended (OEQA) questions. MCQA provides candidate answers along with each question for models to select the correct one (Jang et al., 2017; Xiao et al., 2021). It is favored in answering inference- and explanatory-type questions (e.g., “why and how”) that require relatively longer answers. OEQA treats each answer as a category and requires the models to predict a correct answer from a predefined answer list (Xu et al., 2017; Yu et al., 2019). It is mainly adopted to assess factoid QA where the questions pertain to visual recognition (e.g., “what is”). We refer to the recent survey (Zhong et al., 2022) for a more detailed introduction.
2.2 VideoQA Technique Evolution
The VideoQA techniques generally follow the corresponding development of vision and text representation models. The development can be roughly divided into three stages. The first is 1) CNN + RNN (Jang et al., 2017; Xu et al., 2017; Le et al., 2020; Li et al., 2022; Xiao et al., 2022a) where video and text are encoded by convolutional neural networks (e.g., ResNet (He et al., 2016)) and recurrent neural networks (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) respectively. The second stage 2) CNN/ViT+BERT (Sun et al., 2019; Li et al., 2020; Seo et al., 2021; Lei et al., 2021; Yang et al., 2021; Xiao et al., 2022b, 2023; Fu et al., 2023; Li et al., 2023f, e), features improved visual (Dosovitskiy et al., 2021) and language representations (Devlin et al., 2018; Liu et al., 2019), where the models also benefit from self-supervised cross-modal pre-training at scale and fine-tuned small language models (e.g., BERT and RoBERTa) on the target datasets. Finally, the third (and the current) stage 3) CLIP+LLM (Yang et al., 2022; Liu et al., 2023; Maaz et al., 2023; Li et al., 2023a; Yu et al., 2023; Ko et al., 2023; Zhang et al., 2023b; Dai et al., 2023; Zhang et al., 2023c; Lin et al., 2023) features pre-trained cross-modal visual encoders (e.g., CLIP-ViT (Radford et al., 2021)) and frozen LLMs (e.g., Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023)) with instruction tuning of projection and adaptation modules. In the transition of each stage, there are significant performance jumps. Especially in the LLM stage, leading models like GPT-4V/o (OpenAI, 2023) and Gemini (Team et al., 2023) rival human performance in answering standard visual questions.
2.3 LLMs for VideoQA
Existing LLM-based VideoQA methods can be roughly classified into 3 groups: 1) General-purpose MLLMs, where models are not specialized to VideoQA and are instead capable of handling various vision-language tasks such as image/video QA and captioning. Typical models are InstructBLIP (Dai et al., 2023), Video-ChatGPT (Maaz et al., 2023), Video-LLaMA (Zhang et al., 2023b), VideoChat (Li et al., 2023b), VideoChat2 (Li et al., 2024a), Video-LLaVA (Lin et al., 2023), LLaMA-VID (Li et al., 2024c), and LLaVA-NeXT (Zhang et al., 2024e), and commercial models such as GPT-4V/4o (OpenAI, 2023), Gemini (Team et al., 2023). These models do not finetune on the target datasets but are instruction tuned and prompted to perform different tasks, e.g. VideoQA. We refer to the recent survey (Tang et al., 2023) for a more detailed introduction.
2) Specialized Video-LLMs exploit frozen LLMs or general-purpose image MLLMs (e.g., BLIP-2 (Li et al., 2023a), LLaVA (Liu et al., 2023), LLaMA-Adapter (Zhang et al., 2024b)) while finetuning adaptor or LoRA (Hu et al., 2021) modules on target VideoQA datasets. Successful examples include FrozenBiLM (Yang et al., 2022), SeViLA (Yu et al., 2023), LLaMA-VQA (Ko et al., 2023). Additionally, Video-LLaMA (Zhang et al., 2023b) is a general-purpose Video-LLM, but in this paper we study its finetuned version (Xiao et al., 2024) that is similar to LLaMA-Adapter but with Video Q-Former to encapsulate short clips of a video.
3) Tool-based Video-LLMs. These methods are parameter-free and focused on the use or collaboration of different MLLMs (LLaVA, GPT-4V, LaViLA (Zhao et al., 2023)) and LLMs (LLaMA, GPT-3.5 and GPT-4) as tools (e.g., planning, captioning, and question-answering) for VideoQA. Representative examples include Socratic Models (Zeng et al., 2023), ViperGPT (Surís et al., 2023), MoReVQA (Min et al., 2024), LLoVi (Zhang et al., 2023a), IG-VLM (Kim et al., 2024), TraveLER (Shang et al., 2024), VideoAgent (Wang et al., 2024a; Fan et al., 2024) and VideoTree (Wang et al., 2024b). These methods are training-free and perform zero-shot VideoQA.
2.4 Empirical VQA Study
Existing literature analyzed image VQA with non-LLM techniques (e.g., RNNs or BERT) and focus on specific issues such as data imbalance (Goyal et al., 2017; Kervadec et al., 2021), robustness (Shah et al., 2019), or language bias (Niu et al., 2021). While (Agrawal et al., 2016) analyze the behavior of VQA models from multiple aspects of generalization to long-tailed data as well as complete image and question understanding, they focus on image QA and the techniques examined are from the CNN+RNN stage. In this paper, we will revisit some of the probes (specified in Sec. 3.7) and check if the problems get alleviated in the LLM era.
To our best knowledge, no previous work has comprehensively studied Video-LLMs’ behavior in VideoQA. Several works explore specific phenomena with models from the stage of CNN/ViT+BERT. For example, (Buch et al., 2022) and (Lei et al., 2023) have revealed the single-frame bias issue, where a large number of questions can be answered with a single video frame. (Bagad et al., 2023) have uncovered that existing VLMs (to be distinguished from MLLMs) that fully finetune BERT lack a sense of time. (Zhang et al., 2023d, 2024c) have investigated the problem of imbalanced distribution of candidate answers in MCQA, and show that the models tend to exploit the short-cuts in candidate answers for answer prediction. Recently, (Xiao et al., 2024) have studied visually grounded QA and disclose that existing VLMs and MLLMs are poor at substantiating their predictions with visual evidence.
The above works each focuses on a particular aspect of VideoQA; they study either MCQA or OEQA with non-LLM models. Furthermore, most of they only mention the problems to induce their own solutions or datasets, without delving into the extent of the problems. Given the emerging and revolutionary effect of LLMs for question answering, in this paper, we concentrate on LLM-based techniques and present the first comprehensive analysis for Video-LLMs in VideoQA. We study both MCQA and OEQA either for classification or generation, and additionally compare with non-LLM techniques in face of the same probes.
Additionally, we notice that there is a surge of new benchmarks targeting the analysis of Video-LLMs in a broad aspect of video understanding (Li et al., 2023d, 2024a; Fu et al., 2024; Liu et al., 2024b; Li et al., 2024b) (e.g., diverse video categories such as TV/Magic/Fashion shows and related challenges). In this study, we focus on VideoQA of daily activities for its significance towards human-machine interaction. While we use common names for the probes (e.g., temporal understanding, multimodal reasoning, ), the specific probes and data are entirely different.
| Models | Backbone | Video Encoder | Text Encoder | |
| S-LLM | FrozenBiLM (Yang et al., 2022) | - | CLIP (Radford et al., 2021) | DeBERTa-V2-XL (1B) (He et al., 2020) |
| FrozenGQA (Xiao et al., 2024) | FrozenBiLM | CLIP (Radford et al., 2021) | DeBERTa-V2-XL (1B) (He et al., 2020) | |
| SeViLA (Yu et al., 2023) | BLIP-2 (Li et al., 2023a) | CLIP (Radford et al., 2021) | Flan-T5-XL (3B) (Chung et al., 2022) | |
| LLaMA-VQA (Ko et al., 2023) | LLaMA-Adapter (Zhang et al., 2023c) | CLIP (Radford et al., 2021) | LLaMA (7B) (Touvron et al., 2023) | |
| G-LLM | Video-LLaMA (Zhang et al., 2023b) | BLIP-2 (Li et al., 2023a) | EVA-CLIP (Sun et al., 2023) | LLaMA (7B) (Touvron et al., 2023) |
| VideoChat2 (Li et al., 2024a) | BLIP-2 (Li et al., 2023a) | UMT-L (Li et al., 2023c) | Vicuna (7B) (Chiang et al., 2023) | |
| Video-LLaVA (Lin et al., 2023) | LLaVA (Liu et al., 2023) | LangBind (Zhu et al., 2023) | Vicuna (7B) (Chiang et al., 2023) | |
| LLaMA-VID (Li et al., 2024c) | - | CLIP (Radford et al., 2021) | Vicuna (7B, 13B) (Chiang et al., 2023) | |
| GPT-4o | - | unknown | GPT-4 | |
| T-LLM | LLoVi (Zhang et al., 2023a) | LLaVA, GPT-4 | LangBind (Zhu et al., 2023) | GPT-4 |
| N-LLM | VIOLETv2 (Fu et al., 2023) | VIOLET (Fu et al., 2021) | VidSwin (Liu et al., 2022) | BERT (110M) (Devlin et al., 2018) |
| CoVGT (Xiao et al., 2023) | VGT (Xiao et al., 2022b) | ResNet (He et al., 2016) | RoBERTa (125M) (Liu et al., 2019) | |
| Temp[CLIP] (Xiao et al., 2024) | - | CLIP (Radford et al., 2021) | RoBERTa (125M) (Liu et al., 2019) |
| QA Tasks | Datasets | Video Length | # Answers | # Videos | # QA pairs |
| MCQA | NExT-QA (Xiao et al., 2021) | 42s | 5 | 570 | 5K |
| NExT-GQA (Xiao et al., 2024) | 40s | 5 | 990 | 5.5K | |
| EgoSchema (Mangalam et al., 2023) | 180s | 5 | 500 | 0.5K | |
| Perception Test (Pătrăucean et al., 2023) | 23s | 3 | 5.9K | 19K | |
[…]
3 Probes and Analyses
3.1 Overview
We explore the following probes to examine Video-LLMs’ behavior in VideoQA. The probes are related to either video understanding or question answering:
1.
Temporal Understanding. VideoQA characterizes temporal dynamics compared with image VQA. Hence, our first study is to check if Video-LLMs can (or to what extent) interpret the chronological order of video content by answering paired adversarial temporal questions, i.e., questions elicited by different time words but are about the same video content, e.g., from “A after B” to “B before A”.
2.
Visual Grounding. To discern possible reasons for model failures in following the video content to answer questions, we further study the problem of visual grounding, which is to check whether or to what extent the Video-LLMs’ predictions anchored on the relevant video content versus irrelevant contexts.
3.
Multimodal VQA Reasoning. Both temporal understanding and visual grounding are evaluated on MCQA because there is no suitable OEQA benchmarks for such tests. As such, we delve deeper into multimodal reasoning to check how much Video-LLMs can reason faithfully from video and question to the correct answer, without influence from short-cuts of candidate answers.
4.
Robustness. Are Video-LLMs robust (like human) to general data perturbation? We further study if model decisions are invariant and equivariant to more general video and question perturbations that are not specific to a particular short-cut problem.
5.
Generalizability. All the above experiments curate new data for testing, which may have probed model generalization to a certain extent. Yet for a better study, we consider two classic generalization problems of long-tailed and out-of-distribution (OOD). Long-tailed is to predict long-tailed answers within the same dataset. OOD is to answer questions of completely different types or related to different kinds of videos (e.g., short to long videos, 3rd- to 1st-person view videos) across different datasets.
3.2 Tested Models
We select prominent Video-LLMs that have reported state-of-the-art performance on VideoQA and summarize them in Tab. 1. These models exploit cross-modal pretrained visual encoders (e.g., CLIP (Radford et al., 2021), EVA-CLIP (Sun et al., 2023) and UMT-L (Li et al., 2023c)) and powerful LLMs (e.g., DeBERTa-V2 (He et al., 2020), Flan T5 (Chung et al., 2022), LLaMA (Touvron et al., 2023), Vicuna (Chiang et al., 2023), GPT-4 (OpenAI, 2023)) to perform VideoQA.
Models in the 1st block of Tab. 1 are specially finetuned on target datasets to achieve VideoQA. Models in the 2nd block are general-purpose Video-LLMs and most of they do a zero-shot evaluation on VideoQA. As exception, VideoChat2 (Li et al., 2023b) includes the training data of NExT-QA (Xiao et al., 2021) (one of our major testbed) for instruction tuning, and thus should be considered a normal evaluation versus zero-shot. Also, for Video-LLaMA, we adopt the implementation in (Xiao et al., 2024), which is fine-tuned on NExT-QA by adding adpation parameters as in LLaMA-Adapter (Zhang et al., 2023c). Additionally, we include the recent GPT-4o (OpenAI, 2023) as a representative of commercial models for off-the-shelf use. In the 3rd block, we include LLoVi (Zhang et al., 2023a) which is a tool-based Video-LLM established on LLaVA (Liu et al., 2023), LLaViLA (Zhao et al., 2023) and GPT-4. As non-LLM baselines (the bottom block), we consider VIOLETv2, CoVGT and Temp[CLIP] as representative methods for their good performance achieved by fully finetuning small language models, such as BERT and RoBERTa.
All models except for Video-LLaMA, we use their official checkpoints (or API for GPT-4o). It is noteworthy that different model often tackles a different type of QA, either MCQA or OEQA. For GPT-4o, we uniformly sample video frames and feed them along with the questions into the API and prompt it for QA. Related details are presented in the Appendix .4.
3.3 Datasets & Evaluation Measures
Our major experiments are conducted on NExT-QA (Xiao et al., 2021) for multi-choice QA, and MSVD-QA (Xu et al., 2017) for open-ended QA. We select the datasets for their high quality and popularity with completed models of both non-LLMs and LLMs. Moreover, we add other datasets for specific probes as necessary. Concretely, we test on NExT-GQA (Xiao et al., 2024) for visually grounded QA (Sec. 3.5), NExT-OOD (Zhang et al., 2024c) for multi-choice biases in multimodal VQA reasoning (Sec. 3.6) as well as generalization (Sec. 3.8.2), ActivityNet-QA (Yu et al., 2019) for robustness in answering open-ended temporal questions with shuffled frames (Sec. 3.7). To probe generalization across different types of videos, we add EgoSchema (Mangalam et al., 2023) and Perception Test (Pătrăucean et al., 2023) datasets (Sec. 3.8.2). The former features long-form video understanding from an ego-centric viewpoint, while the latter emphasizes visual perception and reasoning of fine-grained human-object interactions.
Other details are listed in Tab. 2. NExT-OOD shares the same statistics as NExT-QA so we do not list it in the table. Finally, for evaluation, we report both accuracies and flip rates of specific predictions before and after the probes, where the flip rate denotes the proportion of changed predictions.
3.4 Temporal Understanding
VideoQA distinguishes from image VQA for its factoring a rich set of temporal questions signaled by time words “before/after/when”. Video-LLMs demonstrate increasingly high accuracy in answering temporal questions (Xiao et al., 2021), so we try to verify if these models really interpret content order in videos. To this end, we derive two new sets of questions from the original temporal questions in NExT-QA.
Considering the example in Fig. 3. The original question asks the person’s next action after “taking the black item away” and should be answered with “pat the animal”. Our first edit, called Temporal Exchange (TE), exchanges the actions in the question and the correct answer, by asking for the action before “the person pats the animal” (right of Fig. 3). The models should be able to answer “take the black item away” for the derived question if it correctly answered the original question. A second edit called Temporal Description (TD), focuses on the time order and isolates the effect of language correlation by asking in a coarse manner of “what did xx do or what is happening/doing”. The bottom of Fig. 3 shows a TD example, where the question provides little to no contextual information compared to the original and TE questions. Moreover, all answers for the question describe the same visual content with different orders though only one is correct. Correctly answering TD questions requires stronger action order reasoning compared to the original and TE questions.
To obtain the testing examples, we collect about temporal questions (TO) originated from NExT-QA and derive the TE and TD examples by prompting GPT-4 and parsing the syntactic structure of the questions respectively. Regarding prompting for TE examples, we split the questions into two groups of “before/after” and “when”, and design different prompts for them to achieve better results. Regarding parsing for TD questions, we mainly detect the time words “before/after/when”, keep the question parts (e.g., “what did the person do?”) as new questions, and recombine the contextual parts (e.g., “she took the black item away”) with the correct answers (e.g., “pat animal”) while varying the time words to form different options (see Fig. 3). More details are presented in Appendix .1. For both the generated TE and TD examples, we further manually check and refine as necessary to ensure the quality of the derived QAs.
The results in Fig. 2 demonstrate that most models are brittle towards such temporal tests. Specifically, the high flip rates in Fig. LABEL:fig:temp-flip-res show that both LLM and non-LLM methods tend to change their answers for the edited questions. Moreover, the accuracy in Fig. LABEL:fig:temp-acc-res also show that most models’ performances shrink. The effects are specially prominent for the TD test, where most models flip more than 70% of their predictions and decline their accuracy by more than 30%, even though the involved video contents remain the same (also refer to prediction examples in Fig. 4).
Among the models, GPT-4o significantly outperforms others in the TD test even though it does not encode any temporal information (we specify in the prompts that the images are ordered video frames). By jointly considering the accuracy in Fig. LABEL:fig:temp-acc-res, we speculate that the models trained on the target datasets (all models except for GPT-4o and LLoVi (Zhang et al., 2023a)) likely overfit to dataset statistics rather than reason about time faithfully. Additionally, LLaMA-VQA (Ko et al., 2023) shows the smallest flip rate in TE test. We attribute such strength of LLaMA-VQA to its auxiliary training objective by inversely predicting the questions given the answers, which implicitly suits the TE probes by enhancing the relation between the answer and the contextual term in the question. Moreover, Fig. LABEL:fig:temp-acc-res shows that models using advanced LLMs such as LLaMA (Li et al., 2024c) and Vicuna (Chiang et al., 2023) perform better on the TE test but worse on the TD test than non-LLM method CoVGT. This indicates that LLM-based methods are better at exploiting language priors or commonsense in the TE questions for predicting answers, even though they seem weaker than non-LLM methods in understanding video content order. Again, we find that GPT-4o achieves the highest accuracy in both TE and TD tests, showcasing strong zero-shot temporal QA capability.
The above observations, together with the findings in some related studies (Bagad et al., 2023; Liu et al., 2024b), suggest that the end-to-end learning approach (even with LLMs) may not be a good choice to cope with temporal relations. It seems more favourable to explicitly reason across the sequence of video frames in a way like GPT-4o, and alternatively, to convert the video frames into time-aware descriptions and then feed them into LLMs for question answering, such as in MoReVQA (Min et al., 2024) and LLoVi (Zhang et al., 2023a). Yet how to effectively and efficiently coordinate visual captioning and question-answering is still open challenge that deserves further efforts.
3.5 Visual Grounding
With the observation that model answers flip a lot even with paired questions about the same video content, a natural question arises – to what extent are the models’ predictions grounded on the relevant video content? To study this, we conduct experiments on visual grounding. Specifically, we experiment on the recent NExT-GQA dataset (Xiao et al., 2024). This dataset extends NExT-QA with temporal location labels of question-answer pairs. With NExT-GQA, we compare model performance with different visual inputs as follows: NormalVQA follows the common practice of performing VideoQA by uniformly sampling frames over the whole video. BlindQA removes the video inputs to train and test the models with only the QA pairs to check reliance on pure language priors. PosVQA and NegVQA provides only QA-relevant (positive) or only irrelevant segments (negative) as inputs to the model (see Fig. 6 as an illustration). Finally, we evaluate Grounded QA (GQA) accuracy, i.e., jointly answer the questions and also localize the QA-relevant video moments. By comparing QA performance with and without requirements on visual grounding, we study how much the models’ predictions are anchored on the relevant video content. Note that similar experiments are also conducted in NExT-GQA (Xiao et al., 2024), but their analyses are focused on the connection between visual grounding and VQA, with also their major findings established on non-LLM techniques. In this paper, we inherent such settings but target a thorough analysis of Video-LLMs’ behavior in visual grounding and a comparison with non-LLM techniques.
The results in Fig. 5 deliver the following major observations: 1) Fig. LABEL:fig:ground_nbp shows that both LLM and non-LLM methods are poor at substantiating their answers with relevant video content. For example, from NormalVQA to Grounded VQA, i.e. when the correct predictions are required to be visually grounded, the accuracies decrease by more than 40% for all models.
2) Compared with the non-LLM method (CoVGT), specialized Video-LLMs (e.g., FrozenGQA and SeViLA) benefit more from short-cut learning. This is because they remarkably surpass the non-LLM method on normal VQA accuracy (by 10%) but are only marginally better on GQA (e.g., by 2%) (Fig. LABEL:fig:ground_nbp). Also their LLMs alone (e.g., DeBERTa-V2-XL and Flan-T5-XL) already wins over CoVGT’s language model (RoBERTa) by more than 7% according to the corresponding BlindQA accuracies. A further comparison of NormalVQA vs. BlindQA (Fig. LABEL:fig:ground_nbp), NormalVQA vs. PosVQA vs. NegVQA (Fig. LABEL:fig:ground_nnp) indicates that the short-cuts are primarily from language priors and spurious vision-text correlations. For example, the BlindQA baselines without any visual inputs retain more than 70% of the corresponding NormalVQA’ performance. Moreover, visual signals that come from answer-irrelevant video segments yield surprisingly high performance that is competitive to model variants with answer-relevant (ground-truth) visual segments (NegVQA vs. PosVQA).
3) Comparing among models, Fig. LABEL:fig:ground_nbp manifests that the predictions of LLoVi (using GPT-4) and GPT-4o are more visually grounded. Because they achieve the highest GQA accuracy despite being zero-shot methods. Nonetheless, Fig. LABEL:fig:ground_nnp shows that SeViLA and GPT-4o’ predictions are more responsive to different video inputs. Because they have the largest performance gaps between QA with the relevant (positive) and irrelevant (negative) temporal segments. By comparing the BlindQA performance of the finetuned models (CoVGT, FrozenGQA and SeViLA) and zero-shot models (LLoVi and GPT-4o) in Fig. LABEL:fig:ground_nbp, we can conclude that finetuned models are much better at exploiting statistic language biases for QA. For example, a finetuned RoBERTa (CoVGT) performs even better than zero-shot GPT-4, 49.3% vs. 45.3%.
Additionally, in our implementation, when GPT-4o is prompted to also output the indexes of the video frames that it uses to predict the answers versus just answering the questions, its QA accuracy increases remarkably by more than 15%, from 49.0% to 64.7% for NormalVQA. This suggests GPT-4o’s great potential in reasoning like human and providing visually grounded answers if appropriately prompted. We believe this is also an interesting finding and raise it (how to use/prompt MLLMs like GPT-4o) as an open challenge for future exploration.
3.6 Multimodal VQA Reasoning
The analyses of both temporal understanding and visual grounding are based on MCQA because current OEQA benchmarks do not support such tests, e.g., no sufficient temporal questions or no temporal location labels. Thus, we want to study if (or to what extent) the models’ failure of reasoning is because of the short-cuts in candidate answers. We call this probe as multimodal reasoning because such short-cuts dispense with the reasoning between videos and questions, leading to a direct jump to the answers.
Accordingly, we study two kinds of short-cuts related to candidate answers. A video-answer (VA) short-cut exists when only one candidate answer contains the visual objects presented in the video, while most other answers do not involve video-related objects. As such, models may ignore the question and choose the video-relevant answer as final prediction. A similar question-answer (QA) short-cut exists when only one candidate answer contains question-relevant keywords. This shortcut allows models ignore the video and also bypass interpreting the question as well; they may simply associate common keywords in question and answer for prediction. Fig. 7 illustrates such VA and QA short-cuts. Note that the short-cuts may not necessarily bring correct answers but can result in false positives, if for example the video-object related answers are not the correct ones. This is possible because the correct answers often contain the actions (or verbs) presented in the videos though without the objects.
To study the problem, we randomly sample 30% (1.2K to save API budgets) of the validation data from NExT-QA. We then check for the samples that have the aforementioned VA and QA short-cuts. For the target samples, we curate new answers for each question on the basis of the original candidate answers to avoid the possibility of having such shortcuts, i.e., do little change to make most of the candidate answers contain the corresponding video objects or question keywords as shown in the bottom of Fig. 7. In our implementation, we prompt GPT-4 to achieve the goal, followed by human checking and correction as necessary.
Specifically, to eliminate VA short-cut, we first obtain from the related video a collection of visual object labels. We begin by detecting object labels from the sampled video frames by using image-tagging model RAM (Zhang et al., 2024d) and then filter for redundancy before merging the labels together. With the object labels (Examples in the top of Fig. 7) and the original QAs, we then prompt GPT-4 (OpenAI, 2023) to incorporate into each candidate answer the appropriate visual label to make it relate to the video (Examples in the bottom left of Fig. 7).
For QA short-cut, we follow a similar practice to prompt GPT-4 to reword each answer option to contain keywords in the question. Note that the identification of samples with short-cuts are completed at the same time by asking GPT-4 to first check before modifying the answers (details in Appendix .2) . We removed samples whose candidate answers are the same as the original ones as these samples are identified by the GPT-4 as qualified samples without VA and QA short-cuts. Eventually, we obtain two different sets of samples with both size of for probing of VA and QA short-cuts respectively. The high short-cut rates ( out of ) is because our short-cuts are defined on nouns while NExT-QA features verbs in its answers.
The results in Fig. 8 show that almost all models, regardless of the language models used, suffer substantial declines in accuracy (10% 20%) when answering questions with edited distractor answers (see prediction examples in Fig. 9). The flip rates are also high with 21% 37% even though the triplets of video, question, correct answer remain the same. This posits that the models largely rely on the short-cuts in candidate answers and are limited in faithful multimodal reasoning from video and question to correct answers. Again, GPT-4o’s zero-shot performance surpasses others on the edited data.
Compared with non-LLM method (CoVGT), all Video-LLMs, especially VideoChat2, show higher flip rates when confronted with the edited options. This suggests that Video-LLMs are better at exploiting the short-cuts of candidate answers for question answering. By comparing the performances on QA and VA tests, we find that most Video-LLMs experience greater performance loss on edited VA data compared to QA data, especially for GPT-4o whose performance degenerates marginally on QA questions (2.6%) but more on VA type questions (9.9%). Such a discrepancy suggests that the models are primarily exploiting spurious vision-text correspondence for answer prediction.
Additionally, we extend our experiments to the recent NExT-OOD dataset (Zhang et al., 2024c). NExT-OOD is also designed to reduce candidate answer biases in NExT-QA. It aims for a balanced distribution of objects and actions in the correct and distractor answers, whereas our VA and QA short-cuts pertains to the video- and question- correlated candidate answers. The results in Tab. 3 show that Video-LLMs perform surprisingly well on NExT-OOD, surpassing the previous non-LLM SoTA by significant margins. Moreover, the accuracies of Video-LLMs only slightly decrease from NExT-QA to NExT-OOD. We speculate that NExT-OOD unintentionally introduces new and simple patterns that LLMs can exploit, such as mismatches between questions and answers as well as other language issues. This additionally reveals that OOD-data for conventional VideoQA techniques may not be effective OOD for LLM-based methods.
3.7 Robustness
The brittle behavior of models in the above studies prompts us to conduct a more general test on model robustness towards common perturbations on the videos and questions. To this end, we consider two kinds of perturbation: data diversifying and data poisoning, and experiment on both MCQA and OEQA.
For data diversifying, we modify the original questions by rephrasing them or prepending spoken phrases (see examples in the bottom of Fig. 10), both are achieved by prompting GPT-4 with human correction as necessary. Robust models should be invariant and not flip their predictions for this perturbation, since the meanings of questions keep unchanged. For data poisoning, we attack the input data by shuffling video frames, shuffling question words, and removing keywords from the questions (e.g., question words ‘6W1H’222Who, What, When, Where, Why, Which, How, nouns, and verbs). Examples are presented in the first three cases of Fig. 10. Intuitively, models should suffer significant performance loss with poisoned data. For data shuffling, we repeat the experiment 3 times with different random seeds and find there is negligible variance in performance ( 0.1%). Therefore, we report only the median results for brevity. Additionally, for data poisoning, we revisit the issue of “answering incomplete questions” explored by (Agrawal et al., 2016), analogous to human behavior of answering by listening to only a part of the question. We realize this by progressively truncating the question, starting from the first word and increasing the omitted portion.
For probing of model robustness in data diversifying, the results (both accuracy and keep rate (1 - flip rate)) in Fig. LABEL:fig:robust-iv-sp and LABEL:fig:robust-iv-rq show that almost all models deteriorate performances, especially in OEQA. Yet, we surprisingly find that GPT-4o’ accuracy gets improved by 3% when adding spoken phrases, we attribute this strength of GPT-4o to its learning from numerous real-life user cases. Additionally, we highlight the weak behavior of the general-purpose Video-LLMs in answering questions with spoken prefixes in OEQA. For example, the accuracy degradation rates and flip dates of Video-LLaVA and LLaMA-VID are both higher than those of non-LLM method VIOLETv2. Finally, comparing the same model (LLaMA-VID) of different LLM sizes (7B and 13B) reveals that larger LLMs improve accuracy but not necessarily invariance to data diversity.
For better interpretation, we visualize some predictions in Fig. 12. The general purpose Video-LLMs (e.g., LLaMA-VID and Video-LLaVA) tend to change their answers with the edited questions even though the questions’ meanings remain the same. Generally, they prefer giving more specific answers, but always result in hallucinations or ignore the actual questions (e.g., refer to examples at the bottom of Fig. 12). A likely reason could be that the models are not instruction-tuned to answer such edited questions. Another cause could be that the modified questions are often longer and appear more complex than the original ones. Consequently, LLMs are induced for more detailed responses point by point, which increases the risk of hallucination. In conclusion, the experiments demonstrate that the models are not sufficiently robust to language diversifying.
For probing of model robustness in data poisoning, Fig. LABEL:fig:robust-ev-sff shows that after shuffling the video frames, all models’ accuracies have little to no change, be it on MCQA or OEQA. Also, the flip rates are small. The exception, however, is GPT-4o, which has a 4 to 5 higher flip rate than other models. This result echos the findings in Sec. 3.4 and 3.5 where GPT-4o conditions better on the video content for question answering. Yet, for all models, the negligible fluctuation of accuracy indicates a strong static bias in making predictions. To be more precise, we additionally evaluate the effect of frame shuffling on a subset of temporal questions whose answers are related to content ordering, and intuitively, should be more sensitive to frame shuffling. The results in Fig. 13 show that shuffling frames again affects little the model performances in answering even temporal questions. This reflects that the models are stubborn and are not anchored on content order for answering temporal questions. Combining with our previous observation on BlindQA experiments (see Sec. 3.5 and Fig. 5), we conclude that existing Video-LLMs do not really capture video temporality though they ground on videos for answer predictions.
Compared with shuffling frames, Fig. LABEL:fig:robust-ev-sfq shows that model performances degenerate more significantly when shuffling question words. Yet, they still maintain 86%97% of their original performances on MCQA, and 70%80% on OEQA. Comparing among models, Video-LLMs generally show higher accuracy and lower flip rates than non-LLM methods. In particular, the performance of GPT-4o on NExT-QA only decreases marginally by 1.8%. The results indicate Video-LLMs’ superiority in resilience of question poisoning. Nonetheless, this kinds of resilience departs from human intuition in answering questions without syntax.
Additionally, the probes of shuffling frames and questions, combined with the previous temporal probe in Sec. 3.4, suggest that existing models have little to no capability of capturing content position or ordering information in neither language nor video inputs. A major reason could be the weak position embeddings. No matter the entangled absolute position embeddings as in BERT, or the decoupled relative position embeddings in other language models, such embeddings currently do not fully reflect the ordering information and point to a potential direction for further investigation.
For removing question key words, Fig. LABEL:fig:robust-ev-rw shows that models suffer more performance loss and flip more predictions when removing nouns and verbs compared to removing the question words. This holds true for both MCQA and OEQA. The observation differs from (Agrawal et al., 2016) where question words are more crucial than verbs. However, this is reasonable as we analyze VideoQA (vs.image VQA), where the QAs are mostly related to actions.
Aside from removing particular words, Fig. LABEL:fig:robust-truncate revisits the models’ behavior with truncated question inputs as in (Agrawal et al., 2016), but in the context of LLMs. Overall, Video-LLMs show stronger performance in answering incomplete question than non-LLM methods, but the specific behavior slightly varies among MCQA and OEQA. For better analysis and comparison, we follow (Agrawal et al., 2016) to highlight the models’ behavior at half (50%) input length.
Concretely, in MCQA, the keep rate (1 - flip rate) at this half point indicates that the Video-LLMs predict a fixed answer for approximately 75% (refer to SeViLA on NExT-QA) of the samples by listening to merely half the questions. This fixed answer ratio is higher than that of 60% for non-LLM method (refer to CoVGT on NExT-QA), the corresponding accuracy is also higher of 60% (vs. 50%). This suggests that with incomplete question inputs, video-LLMs consistently predict correct answers more so than non-LLM methods. In OEQA, the fixed answer rate of Video-LLM (FrozenBiLM) at the half point is about 40%, which is smaller than that of non-LLM method VIOLET2, e.g., 55%. Yet, a joint consideration of the lower accuracy (30%) suggests that FrozenBiLM converges less to the wrong answers. Together with the observations of (Agrawal et al., 2016) for CNN+RNN and attention models, it seems that models with better accuracy are less likely to predict wrong answers even with incomplete question inputs.
Additionally, Video-LLMs achieve higher MCQA accuracy (e.g., 60% of SeViLA) even without the questions. This result echoes our findings in Sec. 3.6 on video-answer short-cuts for decision making. Regarding OEQA, an interesting observation is that VIOLETv2 outperforms FrozenBiLM when fed with incomplete questions, but such superiority gradually disappears with the increase of question words. We suspect this is because VIOLETv2 finetunes on BERT (vs. frozen LLMs) with masked language modelling. The masking mechanism allows cross-modal matching with partial and incomplete sentences, and may conflict with the underlying intent of the probe.
3.8 Generalizability
Video-LLMs, by virtue of using LLMs, are assumed to encode some common sense and open-world knowledge (Kojima et al., 2022; Ko et al., 2023). To that end, their decisions should be less affected by the QA training data and generalize better. Our previous probes implicitly test model generalization to a certain extent by curating new data. This section studies two classic generalization scenarios more explicitly: generalization to long-tailed data within dataset and generalization to out-of-distribution (OOD) data across datasets. Both long tails and OOD data conflict and do not reflect the training data on which the models are developed. It is worth mentioning that the general-purpose and tool-based Video-LLMs have already shown excellent zero-shot performance in our previous probes. Thus, we focus on analyzing the specialized Video-LLMs that are fintuned on the target datasets in this section.
3.8.1 Generalization to long-tailed data
For this experiment, we investigate the models’ predictions at a per-answer basis. We consider open-ended QA with close-set classification setting where each answer is treated as a class label, e.g. MSVD-QA (Xu et al., 2017). Specifically, we analyze an answer’ frequency and its accuracy, as well as the frequency of it being a false positive prediction. An answer’s accuracy stands for the proportion of correctly answered questions over all questions with that answer as ground-truth answer. We experiment with FrozenBiLM and VIOLETv2 as representatives of Video-LLMs and non-LLMs for close-set classification respectively.
We present the results of FrozenBiLM in Fig. LABEL:fig:fair. VIOLETv2 shares almost the same distribution and is omitted from the plot for clarity. The answers’ frequency distribution and accuracy distribution has a correlation coefficient of only for both FrozenBiLM and VIOLETv2. In contrast, the coefficient value between answers’ frequency distribution and the frequency distribution of an answer being a false positive prediction is 0.9 for both methods. The experimental results suggest that there is little to no correlation between an answer’s frequency and its accuracy for both LLM and non-LLM methods. However, there is a strong positive correlation between an answer’s frequency and the frequency of it being a false positive prediction, and this stands true for both LLM- and non-LLM based methods. While we suspect that the LLM in FrozenBiLM may not be large enough (1B), the experiments indeed suggest that finetuned (or specialized) Video-LLMs are similar to non-LLM method in generalizing to long-tailed data distribution.
3.8.2 Generalization to OOD data
The test samples in all previous probes are either derived from the same types of questions or from the same dataset. Here, we focus on generalization across different question types and datasets.
For different question types, we experiment with NExT-QA (Xiao et al., 2021) as it has three distinct question groups: Causal (“why and how”), Temporal (“what before/after/when”) and Descriptive (“what/where is”). We investigate models’ behavior when training with data from one of the three groups and testing with samples from the other two. For different datasets, we test the models pretrained on NExT-QA on two recent datasets: EgoSchema (Mangalam et al., 2023) for 1st-person view long video understanding and Perception Test (Pătrăucean et al., 2023) for fine-grained visual perception and reasoning. These datasets are selected as they are considered one of the most prominent benchmarks for understanding 1st-person view visual contents and fine-grained human-object interactions, respectively.
Fig. 15 shows that Video-LLMs generally maintain stronger QA performance when transferring across different question types compared to non-LLM method (e.g., CoVGT). A comparison among figures of (a)/(b)/(c) suggests that training on temporal/causal questions often yields better performance than training on simple descriptive questions. In particular, Video-LLMs (SeViLA and LLaMA-VQA) have marginal or no accuracy declines on descriptive questions when trained with temporal questions as opposed to training on the corresponding descriptive questions; performance even improves when training with causal questions. While a possible reason is that there are more causal and temporal questions than descriptive questions in the training data, we intriguingly find that the non-LLM method contrasts with LLM method and achieves the highest accuracy of 67.1% (vs. 58.8% and 56.9%) on descriptive questions when trained with descriptive questions. Such discrepancy thus reflects Video-LLMs’ superior behavior in generalizing to descriptive questions even if they are not trained to do so. Yet, we note that the best results for all model across different question groups are always obtained by learning on the full training set.
Among the models, SeViLA generalize the best among question types. However, its superiority disappears when generalizing to EgoSchema (Mangalam et al., 2023). For example, its accuracy is 32.8% which is even worse than that of non-LLM method Temp[CLIP], e.g. 33.8% (refer to Fig. 16). A prediction example in Fig. 17 suggests that SeViLA fails to holistically understand the long video content. This is reasonable as SeViLA identifies only a limited number of keyframes (4 out of 32) for QA, which is a departure from EgoSchema which defines temporal certificate to challenges long-form video understanding. Nevertheless, all Video-LLMs remain their superior performance over non-LLM method (Temp[CLIP]) in generalizing to Perception Test (Pătrăucean et al., 2023). The observations, together with our experiments on NExT-OOD in Sec. 3.6 and the experiments in (Mangalam et al., 2023), suggest that Video-LLMs, even the specialized or finetuned ones, commonly generalize well to OOD data. Yet this is not guaranteed for each specific model.
4 Discussion & Future Directions
4.1 Robustness and Trustworthiness
All our probes showed that the open-sourced Video-LLMs perform well for standard question-answering. However, these models are susceptible to adversarial language modifications while remain stubbornly unaffected to analogous video modifications. These findings reveal a lack of robustness, underscoring the urgent need for interpretability in Video-LLM development. There are many directions for improvement. We outline some of them, dividing them up from the perspective of data, modeling, and learning.
From a data perspective, a direct solution could be to apply our strategies of designing adversarial probes to augment the training data. Our simple attempt (see Appendix .5) shows that this can effectively improve the model performance on the probe data and the original test data. Aside from training data, new video testing benchmarks are emerging (Li et al., 2024a; Fu et al., 2024; Liu et al., 2024b; Li et al., 2024b). Yet almost all of them target comprehensive video understanding, with little regard for interpretability. Related benchmarks such as VIP (Himakunthala et al., 2023) and NExT-GQA (Xiao et al., 2024) are thus valuable efforts in developing trustworthy Video-LLM techniques and deserve more research attention.
From a modeling perspective, VideoQA models are imbalanced in terms of their language and visual representations. The language encoders are drawn from ever-larger and more powerful language models which dominate over comparatively weaker visual representations, where CLIP Radford et al. (2021) and its variants Ilharco et al. (2021); Sun et al. (2023) remain de-facto standards even though they struggle to represent fine-grained details. More importantly, from an image to a video level, the temporal and dynamic modeling tends to be primitive. Hence, exploring and improving the image or video encoders would contribute to more robust models.
Finally, from a learning perspective, existing models learn to directly output the final answers in an auto-regressive manner. Directly targeting the answers (especially lengthy ones) enables the models to ignore the visual contents and take whatever shortcuts they can (e.g., hallucinate based on commonsense knowledge from LLMs) to make plausible but often incorrect responses (Sec. 3.7). As such, we believe the focus on intermediate reasoning steps and post-verifications can benefit model developing towards robustness and trustworthiness. Some recent studies in multimodal chain-of-thought (Wei et al., 2022; Zhang et al., 2023e; Fei et al., 2024) and modular multi-agents (Min et al., 2024; Shang et al., 2024) provide insights in this direction and show improved robustness and interpretability.
4.2 Efficient and Fine-Grained Long-Range Understanding
GPT-4o’s superior behavior in VideoQA (see Fig. 1 and 5), as well as the ascent of SOTA performance on benchmarks like MSVD-QA and MSRVTT-QA suggests that coarse and holistic understanding of short videos (less than 20s) are solvable by scaling up existing Video-LLMs or prompting powerful image MLLMs (like GPT-4o). Yet the understanding of fine-grained details (Xiao et al., 2021) or long-range videos (Mangalam et al., 2023; Fu et al., 2024) remains out of reach. These videos offer additional challenges that require more dedicated analysis and architecture design.
One such hurdle comes from the ‘needle in the haystack’ problem - the search in the long video for short moments (or keyframes) that carry the correct answer. Existing benchmarks such as NExT-GQA (Xiao et al., 2024) and QAEgo4D (Datta et al., 2022) embody such challenges, yet effective techniques are lacking. Recent solutions (e.g., MoReVQA (Min et al., 2024), VideoAgent (Wang et al., 2024a; Fan et al., 2024) and TraveLER (Shang et al., 2024)) show promise. However, they are relatively inefficient, in that they traverse the entire video frame-by-frame, or segment-by-segment, and perform LLM inference multiple times. Efficient understanding of such long videos is gaining attention but continues to be an open problem for the foreseeable future.
The second challenge is a fine-grained understanding of long-range videos. The concept of large temporal certificates, as defined in EgoSchema (Mangalam et al., 2023), highlights accurately differentiating the visually very similar long activities (e.g., teaching someone to play vs. play with someone). Our studies in Sec. 3.8.2 show that existing Video-LLMs targeting short clips do not generalize well to such videos. Traversal-based solutions are ineffective as the task requires reasoning over and weaving together (potentially distant) frames or segments to achieve holistic understanding. Recent memory- and caption-based solutions (e.g., MoReVQA (Min et al., 2024), LLoVi (Zhang et al., 2023a) and VideoAgent (Fan et al., 2024)) as well as the direct long-range encoding solutions (e.g., LongVA (Zhang et al., 2024a)) offer some insights. However, fine-grained understanding, especially in an efficient manner, remains a key challenge.
4.3 Egocentric and Embodied QA
We primarily analyzed VideoQA with exo-centric datasets - videos from the third-person view. It should be noted that an important real-world application for VideoQA is geared towards reliable embodied assistance (Grauman et al., 2022; Majumdar et al., 2024). In this regard, ego-centric, or first-person view videos and associated question-answer pairs are of great interest and relevance. Sec. 3.8.2 shows that Video-LLMs pre-trained on third-person view VideoQA datasets can generalize to these application scenarios. However, their performance is weak compared with the tool-based Video-LLMs (Zhang et al., 2023a) that directly use ego-centric pretrained visual descriptions (Zhao et al., 2023), and well-prompted general-purpose MLLMs (e.g. GPT-4o) (Liu et al., 2024a). Therefore, the efforts of leveraging existing MLLMs for ego-centric VideoQA will be of critical importance in initiating successful embodied assistance applications.
5 Conclusion
This paper has vetted Video-LLMs’ performance in VideoQA by probing their success and failure modes with well-designed adversarial tests. While Video-LLMs generally show better QA accuracy, the decrease rates and flip rates align with or even higher than non-LLM methods when faced with adversarial challenges. Specifically, Video-LLMs show significant limitation in coping with video temporality, of both reasoning the chronological content order and grounding the temporal moments to substantiate the answers. Also, they are unresponsive towards video perturbation while being susceptible to simple language variations of questions and candidate answers. Additionally, Video-LLMs, after finetuned, are not necessarily generalize better. Understanding these limitations is crucial for developing future Video-LLMs and VideoQA techniques, where we also conclude some promising directions to proceed.
Limitations
We note that Video-LLM research is emerging, and there are other powerful models we may have missed testing. We thus release all our test data: https://github.com/doc-doc/VideoQA-LLMs.
Acknowledgements
We greatly thank OpenAI for offering us with research access API credits.